OpenStudy (ujjwal):

Find equivalent polynomial equation of \(\sqrt{a^2-4ab+4b^2}\)

OpenStudy (ujjwal):

*expression

OpenStudy (anonymous):

\[\sqrt{(a-2b)^2}\]\[\left| a-2b \right|\]

OpenStudy (ujjwal):

why use modulus value?

OpenStudy (anonymous):

\[x^2=a \]\[x=\pm \sqrt{a}\]so instead of plus minus we write modulus!